# Product of Experts for Visual Generation

### v1

@@ 4 Experiments
We now show results applying the PoE framework to image/video synthesis tasks. Expert configurations can adapt to the desired, task-dependent input granularity, including high-level text instructions (Sections 4.2, 4.3 and 4.4) and precise, low-level pose controls (Sections 4.2 and 4.3), and leverage expert knowledge ranging across natural image and video priors (Sections 4.2, 4.3 and 4.4), precise physics rules from physical simulators (Section 4.3), and semantic visual understandings from VLMs (Section 4.4).
4.1 Implementation Details
As described in Eq. 4, a generative expert can be defined over the entire scene $x\in\mathbb{R}^{d}$ or a region of the scene $x_{i}\in\mathbb{R}^{d^{(i)}}$ with $d^{(i)}\leq d$, i.e., the expert has prior over only part of the scene, e.g., a local spatial region in Section 4.4. In these cases, we zero-pad the velocity prediction from the generative expert from $\mathbb{R}^{d^{(i)}}$ to $\mathbb{R}^{d}$ before velocity composition (Algorithm 1, line 5).
In other cases, one may restrict the influence of a generative expert to update only a scene region, even when the expert is defined over the full scene $x\in\mathbb{R}^{d}$, e.g., when the expert only has knowledge over foreground object informed by the graphics engine or physics simulators in Sections 4.2 and 4.3. We apply a local (spatial for images and spatio-temporal for videos) weight $\lambda^{(i)}(x)\in\left[0,1\right]^{d}$ on the velocity prediction $v_{t}^{(i)}(x)$ in these cases, where $\lambda^{(i)}(x)$ is the Gaussian-blurred foreground mask.
@@ 4.2 Graphics-Engine-Instructed Image Editing
Task Specifications.
We apply our framework on the task of image editing where user provides an image to be edited, a 3D asset posed in a graphics engine, Genesis [^2] https://github.com/Genesis-Embodied-AI/Genesis, to precisely describe where to insert the asset, and a text description describing higher-level information such as object materials (“metal”) and semantics (“coke can”). We deploy two generative experts: a depth-to-image model, FLUX.1 Depth [dev] flux2024 , and an inpainting model, FLUX.1 Fill [dev] flux2024 . The former provides necessary constraints on the outputs to ensure objects inserted follow input pose specifications, and the latter provides a natural image prior to produce realistic visual outputs.
@@
The task is evaluated on a dataset with 10 natural images paired with 3 object inputs, resulting in 30 scenes in total, with examples shown in Fig. 1. Evaluation metrics cover three major visual aspects, including controllability, image quality, and text alignment. Specifically, we evaluate the MSE distance between generation outputs and input images on background pixels to evaluate background preservation, and the LPIPS zhang2018perceptual distance between outputs and graphics engine’s RGB renderings on foreground pixels to measure the fidelity to input objects. We report aesthetic score using the LAION-Aesthetic predictor [^3] https://github.com/LAION-AI/aesthetic-predictor and measure text alignment with ImageReward xu2023imagereward . Finally, for each evaluation aspect, we query GPT-4o for automatic evaluation. See details in Section F.1.
@@ Results.
We compare our method with an inversion-based image editing baseline, RF-Solver wang2024rfsolver , which receives the same input image and a target text prompt, e.g., “a metal box standing on the ground”, which is inherently coarse and therefore does not produce results conforming to the graphics engine inputs, and it sometimes fail to faithfully preserves input background as shown in Fig. 2. Table 1 suggests that our method achieves better controllability on the task.
Ablation on Conditional Sampling.
Section 3.1 introduced the conditionally sampling strategy, which is crucial for efficient sampling from the product distribution. Removing the updates in Eq. 7 (no-cond in Table 1) decreases performance and results in less harmonized visual outputs (Fig. 2).
Figure 3: Application on Physical-Simulation-Instructed Video Generation. Given an input image and an input sequence from physical simulators describing precise object motions, our method generates videos aligned with input object motions while synthesizing natural video content for non-foreground regions.
Figure 4: Baseline Comparisons on Physics-Simulator-Instructed Video Generation. All frames are processed in grayscale and overlaid with tracking trajectories predicted by SpaTracker xiao2024spatialtracker for visualization.
Methods Controllability Video Quality Semantic Alignment
LPIPS ($\downarrow$) GPT-4o ($\uparrow$) Smooth ($\uparrow$) Aesthetic ($\uparrow$) Imaging ($\uparrow$) GPT-4o ($\uparrow$) ViCLIP ($\uparrow$) GPT-4o ($\uparrow$)
Object-Centric Simulation Input
Traj2V $0.109$±$\pm 0.033$ $0.638$±$\pm 0.172$ $\underline{0.993}$±$\pm 0.003$ $0.534$±$\pm 0.071$ $0.570$±$\pm 0.154$ $0.725$±$\pm 0.179$ $0.258$±$\pm 0.041$ $0.757$±$\pm 0.141$
Depth2V $\mathbf{0.103}$±$\pm 0.031$ $\underline{0.650}$±$\pm 0.144$ $0.984$±$\pm 0.015$ $0.534$±$\pm 0.069$ $0.581$±$\pm 0.155$ $0.750$±$\pm 0.126$ $\underline{0.261}$±$\pm 0.041$ $0.775$±$\pm 0.130$
Image2V $0.111$±$\pm 0.021$ $\mathbf{0.708}$±$\pm 0.104$ $0.978$±$\pm 0.025$ $\mathbf{0.553}$±$\pm 0.057$ $\underline{0.611}$±$\pm 0.101$ $\underline{0.775}$±$\pm 0.099$ $0.255$±$\pm 0.036$ $\underline{0.788}$±$\pm 0.079$
Ours $\underline{0.104}$±$\pm 0.030$ $\mathbf{0.708}$±$\pm 0.104$ $\mathbf{0.994}$±$\pm 0.003$ $\underline{0.549}$±$\pm 0.068$ $\mathbf{0.625}$±$\pm 0.117$ $\mathbf{0.842}$±$\pm 0.132$ $\mathbf{0.270}$±$\pm 0.038$ $\mathbf{0.817}$±$\pm 0.080$
Full-Scene Simulation Input (PhysGen3D dataset)
PhysGen3D $0.438$±$\pm 0.099$ $\mathbf{0.995}$±$\pm 0.002$ $0.571$±$\pm 0.114$ $0.673$±$\pm 0.046$ $0.513$±$\pm 0.117$ $0.224$±$\pm 0.026$ $0.588$±$\pm 0.117$
Inversion $0.237$±$\pm 0.084$ $0.263$±$\pm 0.122$ $\underline{0.993}$±$\pm 0.002$ $0.468$±$\pm 0.097$ $0.410$±$\pm 0.144$ $0.250$±$\pm 0.206$ $0.202$±$\pm 0.058$ $0.400$±$\pm 0.206$
Depth2V $\underline{0.164}$±$\pm 0.065$ $\underline{0.550}$±$\pm 0.173$ $\underline{0.993}$±$\pm 0.003$ $\underline{0.579}$±$\pm 0.067$ $\underline{0.680}$±$\pm 0.081$ $\underline{0.662}$±$\pm 0.216$ $\mathbf{0.242}$±$\pm 0.030$ $\underline{0.788}$±$\pm 0.108$
Image2V $0.242$±$\pm 0.077$ $0.525$±$\pm 0.192$ $0.986$±$\pm 0.010$ $\mathbf{0.594}$±$\pm 0.074$ $0.662$±$\pm 0.052$ $0.638$±$\pm 0.245$ $0.236$±$\pm 0.028$ $\underline{0.788}$±$\pm 0.105$
Ours $\mathbf{0.136}$±$\pm 0.047$ $\mathbf{0.587}$±$\pm 0.220$ $\underline{0.993}$±$\pm 0.003$ $0.575$±$\pm 0.071$ $\mathbf{0.688}$±$\pm 0.045$ $\mathbf{0.763}$±$\pm 0.132$ $\underline{0.239}$±$\pm 0.038$ $\mathbf{0.825}$±$\pm 0.141$
Table 2: Physics-Simulator-Instructed Video Generation.
4.3 Physical-Simulator-Instructed Video Generation
Task Specifications.
The setup from Section 4.2 can be directly extended to dynamic scenes for video generation, with object poses and dynamics (e.g., a ball bouncing, Fig. 3 top row) specified by a physical simulator, an input image, and a textual scene description. We use Genesis as the simulator, and use two generative experts: an image-to-video model, Wan 2.1 wan2025 , and a depth-to-video model [^4] We use the released checkpoint on Wan 2.1 from https://github.com/aigc-apps/VideoX-Fun..
Evaluation.
We evaluate our method and several monolithic models, including image-to-video (Image2V), depth-to-video (Depth2V), and trajectory-to-video (Traj2V), all using the same Wan2.1 (14B) model as backbone, on a dataset consisting of 3 simulation inputs paired with 4 input images each, giving a total of 12 scenes with examples in Fig. 3. To obtain the initial frame for all methods, we edit the input image with our method described in Section 4.2 to insert the simulated object. Depth maps required as method inputs are rendered from the physical simulator, and input trajectory maps are computed using object centroids. We report metrics similar to Section 4.2, with additional ones adopted from VBench huang2024vbench : motion smoothness scores introduced in VBench huang2024vbench , MUSIQ ke2021musiq for imaging quality, and ViCLIP wang2023internvid for overall semantic consistency with input text prompts.
In the above setting, simulation inputs only describe foreground objects. We further evaluate on cases where full-scene simulation is available. We use the simulation results from 9 released scenes from PhysGen3D chen2025physgen3d , a recently proposed method that reconstructs and animates images in physical simulators, as evaluation inputs. The full-scene simulation is closely aligned with natural videos, and therefore we use the noise vector from flow ODE inversion for particle initialization as opposed to random noise (Algorithm 1, line 2). We further compare with flow inversion baseline for this setting.
Results.
When only object-centric simulation is available (Table 2, top half), depth-to-video and trajectory-to-video alone fail to capture rich non-foreground-object motions, such as cars moving and birds running in Fig. 4. In particular, they tend to compensate for the object’s downward motion with upward camera motions, as visualized by the tracking trajectories in Fig. 4. With full-scene simulation as inputs (Table 2, bottom half), our method improves the video quality over the input (e.g., the apple’s black regions in Fig. 5) while remaining faithful. On the other hand, the depth-conditioned baseline deviates from the input simulation as shown in Fig. 5. In both cases, the image-to-video model does not adhere to the input simulation due to the coarseness of text descriptions.
Figure 6: Text-to-Image Generation using regional generative experts and VLM as discriminative expert.
4.4 Text-to-Image Generation with Regional Composition
We further demonstrate the efficacy of distribution composition on the task of text-to-image generation, where we first query an LLM to parse an input text prompt into spatial bounding boxes with regional prompts (Fig. 6 left), aiming to compose the distributions from text-to-image model (FLUX.1 [dev]) conditioned on local text prompts, one for each local region, as generative experts, and from VQAScore lin2024evaluating as a discriminative expert, which gives a scalar value predicting the probability that its input image is aligned with the input text prompt. We compare our methods using $L=16,4$ particles and $1$ particle (which amounts to not using the discriminative expert), a variant using no generative composition (equivalent to best-of-N sampling for vanilla FLUX) while matching the compute budget for 16 particles, and with the vanilla FLUX backbone. In the examples shown in Fig. 6, our compositional approach achieves better text alignment compared to the baselines, and the discriminative expert provides crucial knowledge, e.g., counting, to steer the output distribution.
@@ 5 Conclusion
We have introduced a Product of Experts (PoE) framework for composing knowledge from heterogeneous sources at inference time for visual generation, and proposed an efficient sampling algorithm leveraging Sequential Monte Carlo. Our training-free approach effectively integrates visual generative priors, VLM guidance, and physics-based constraints. We demonstrated that this method generates images and videos with higher complexity and improved controllability compared to monolithic approaches, providing a promising direction for the next generation of visual generation models.
Acknowledgments.
This work was in part supported by ONR N00014-23-1-2355, ONR YIP N00014-24-1-2117, ONR MURI N00014-22-1-2740, and NSF RI #2211258. YZ is in part supported by the Stanford Interdisciplinary Graduate Fellowship.
@@ Appendix A Overview
The supplementary material contains discussions (Appendix B), extended descriptions of the framework (Appendices C and D), and additional experimental results (Appendix E) and details (Appendix F).
@@ Appendix B Discussions
All image experiments are run using 1 NVIDIA H200 GPUs and all video experiments with 2 NVIDIA H200 GPUs. Generating one sample in Sections 4.2, 4.3, 4.4 and C takes approximately 4min, 5min, 4min, and 30min, respectively.
@@
The proposed method requires extra computation due to intermediate MCMC steps and parallel sampling compared to vanilla feedforward inference of the image or video generative model backbones. Furthermore, the sampling efficacy relies on the assumption that expert predictions are reasonably compatible, and strong incompatibilities might require other sampling techniques or training. This work proposes a framework and provides empirical validation with small number of experts and we leave large-scale PoE sampling as future work.
@@
Appendix C Framework Instantiation with Autoregressive Experts
As described in Section 3.3, instantiating the framework requires constructing the annealed distributions $\{p_{t}(x)\}_{t=1}^{T}$in Eq. 3, and defining the corresponding transition kernel from $p_{t}(x)$ to $p_{t-1}(x)$. The main paper uses flow ODE discretization to construct the annealed distributions.
An alternative construction is to use autoregressive marginals $p_{t}^{(i)}(x):=\bar{p}^{(i)}(x_{1:t})$ such that $x_{1:T}=x$. Note that we flip the order of $t$ to traverse from $1$ to $T$ for notation convenience. We use bar notations for pre-trained models in the following discussions. Given an autoregressive model that allows for sampling from $\bar{p}(x_{1:t+1}\mid x_{1:t},\theta_{i})$, Algorithm 1 can be implemented with $\mathcal{K}_{t\leftarrow t-1}(x^{\prime}\mid x):=\prod_{i=1}^{N}\bar{p}^{(i)}(x_{1:t}^{\prime}\mid x_{1:t-1})$ and $\mathcal{K}_{t}$ as Gibbs sampling.
The following are empirical results under this construction. We use two generative experts: a frame-wise autoregressive video generation model, FramePack zhang2025framepack , and a graphics engine, Genesis, which defines a Gaussian distribution centered on a video sequence of the physics simulator’s RGB rendering overlayed with the input initial frame, visualized as “Input Sim” in Fig. 7. In this case, $x_{1:t}$ represents the first $t$ frames of a video $x$, and the product distribution is defined as $p(x)\propto e^{-w\|x-c^{\text{simulation}}\|^{2}}p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$ with some constant $w\in\mathbb{R}$. We approximate sampling from this distribution with first sampling from $p^{\text{image2video}}$, then do 8 gradient updates with respect to $\mathcal{L}_{2}$ loss $\|x-c^{\text{simulation}}\|^{2}$.
Qualitative results for the application from Section 4.3 and comparisons with vanilla image-to-video inference using the same backbone model, HunyuanVideo kong2024hunyuanvideo  [^5] Pre-trained weights from https://huggingface.co/lllyasviel/FramePack_F1_I2V_HY_20250503., are shown in Fig. 7, with $81$ frames per sequence and spatial resolution $512\times 768$. We set $L=1$, with 1 MCMC sampling step per iteration for this experiment. Results from our method follow the input physics simulation while synthesizing motions from other subjects in the scenes following the natural video prior from the backbone model.
Figure 7: Framework Instantiation with Autoregressive Annealed Distributions. Given an input first frame (first column), a text prompt (at the bottom of each subfigure), and the RGB rendering from a physics simulator overlayed with the input first frame (top row of each subfigure), our method generates natural videos following the object motion specified by physics simulations. In comparison, the baseline image-to-video model does not produce the desired object motion.
Appendix D Framework Details
D.1 Flow Models
Consider the $i$-th generative expert, which is a flow model that defines a time-dependent continuous probability path $\bar{p}^{(i)}(x;\bar{t}),x\in\mathbb{R}^{d},\bar{t}\in\left[0,1\right]$, with boundary conditions $\bar{p}^{(i)}(x;0)=\mathcal{N}(0,I)$ and $\bar{p}^{(i)}(x;1)\approx p_{\text{data}}^{(i)}(x)$. Each flow model is associated with a scheduler $\bar{\alpha},\bar{\sigma}:\left[0,1\right]\rightarrow\mathbb{R}$ that defines a conditional probability path $\bar{p}^{(i)}(x\mid y;\bar{t})=\mathcal{N}(x\mid\bar{\alpha}(\bar{t})y,\bar{\sigma}(\bar{t})I),y\sim p_{\text{data}}^{(i)}(y)$, and is trained to predict a velocity field $\bar{v}^{(i)}(x;\bar{t},\theta_{i})$ that generates the marginalized path $\bar{p}^{(i)}(x;\bar{t})$, where $\theta_{i}$ denotes model parameters. While the training-time scheduler $(\bar{\alpha},\bar{\sigma})$ can differ across experts, they can be aligned during inference time via a scale-time transformation lipman2024flow .

### v2

@@ 4 Experiments
We apply our framework to image/video synthesis tasks. Expert configurations can adapt to the desired, task-dependent input granularity, including high-level text instructions (Sections 4.1, 4.2 and 4.3) and precise, low-level pose controls (Sections 4.1 and 4.2), and leverage expert knowledge ranging across natural image and video priors (Sections 4.1, 4.2 and 4.3), precise physics rules from physical simulators (Section 4.2), and semantic visual understandings from VLMs (Section 4.3).
@@ 4.1 Graphics-Engine-Instructed Image Editing
Task.
This task provides the following image object insertion interface for applications where users want to insert an object precisely into certain parts of the image. Inputs consist of an image to be edited, an input 3D asset posed in a graphics engine, and a text description describing higher-level information such as object materials (“metal”) and semantics (“coke can”). We deploy two generative experts for this task: a depth-to-image model, FLUX.1 Depth [dev], and an inpainting model, FLUX.1 Fill [dev] (FLUX, 2024). The former ensures inserted objects follow input pose specifications, and the latter provides a natural image prior to produce realistic outputs.
@@
We compare with an editing method, RF-Solver (Wang et al., 2024), which inverts the input image to a Gaussian noise and then generates the output starting from that noise, conditioned on a text prompt that additionally describes where to insert the object, e.g., “a metal box standing on the ground”. The evaluation dataset consists of 10 natural images paired with 3 object assets, resulting in 30 scenes in total, with examples in Fig. 1. The metrics cover three major aspects: controllability, image quality, and text alignment. Specifically, we evaluate the MSE distance between generation outputs and input images on background pixels to evaluate background preservation, and LPIPS (Zhang et al., 2018) distance between outputs and graphics engine’s RGB renderings on foreground pixels to measure the fidelity to input objects. We report aesthetic score using the LAION-Aesthetic predictor (Schuhmann & Beaumont, 2022) and measure text alignment with ImageReward (Xu et al., 2023), and query GPT-4o for automatic evaluation (details in Section E.2).
Results are in Table 1 (top half). Baseline outputs do not conform to input object pose due to the lack of precision of text prompts, and may fail to preserve input background as observed in Fig. 2.
Ablation on Conditional Sampling.
Section 3.1 introduced the conditionally sampling strategy, which is crucial for efficient sampling from the product distribution. Removing the updates in Eq. 6 (“No Cond” in Table 1) decreases performance and results in less harmonized visual outputs (Fig. 2).
Character Insertion.
The task spec above is closely relevant to the character insertion task studied in Magic Insert (Ruiz et al., 2024), where the goal is to insert a character from an image into a background image. We use 10 background images paired with 8 character inputs released on their official demo page to construct an evaluation dataset of 80 scenes. Since their model uses a different backbone (Podell et al., 2023), for fair comparisons, we also include three baselines using the same FLUX backbone as ours: Add-it (Tewel et al., 2024), SDEdit (Meng et al., 2021), and FLUX-Fill.
Results are in Table 1 (bottom half). Add-it receives text instructions for object insertion, but tends not to follow the specified character location in the prompts; SDEdit produces harmonized outputs but does not closely preserve the background due to the global noising operation; the inpainting method, FLUX-Fill, sometimes ignores the character descriptions in the input text instruction and fails to insert any object. These observations are also reflected qualitative samples in Section D.1.
Methods Controllability Video Quality Semantic Alignment
IoU (fg) ($\uparrow$) LPIPS ($\downarrow$) GPT‑4o ($\uparrow$) Smooth ($\uparrow$) Aesthetic ($\uparrow$) Imaging ($\uparrow$) GPT‑4o ($\uparrow$) ViCLIP ($\uparrow$) GPT‑4o ($\uparrow$)
Object-Centric Simulation Input
[…]
Ours $\underline{0.739}$$\pm.221$ $\underline{0.104}$$\pm.030$ $\mathbf{0.708}$$\pm.104$ $\mathbf{0.994}$$\pm.003$ $\underline{0.549}$$\pm.068$ $\mathbf{0.625}$$\pm.117$ $\mathbf{0.842}$$\pm.132$ $\mathbf{0.270}$$\pm.038$ $\mathbf{0.817}$$\pm.080$
Full-Scene Simulation Input (PhysGen3D (Chen et al., 2025) dataset)
[…]
Depth2V $\underline{0.164}$$\pm.065$ $\underline{0.550}$$\pm.173$ $\underline{0.993}$$\pm.003$ $\underline{0.579}$$\pm.067$ $\underline{0.680}$$\pm.081$ $\underline{0.662}$$\pm.216$ $\mathbf{0.242}$$\pm.030$ $\underline{0.788}$$\pm.108$
[…]
Table 2: Physics-Simulator-Instructed Video Generation Evaluation.
Figure 4: Comparisons on Physics-Simulator-Instructed Video Generation. Predictions are processed in grayscale and overlaid with estimated tracking (Xiao et al., 2024) for visualization.
4.2 Physical-Simulator-Instructed Video Generation
Task.
The setup from Section 4.1 can be directly extended to dynamic scenes for video generation, with object poses and dynamics (e.g., a ball bouncing, Fig. 3 top row) specified by a physical simulator, an input image, and a textual scene description. Below, we consider two sets of experts: flow-based (Wang et al., 2025) and autoregressive-based (Zhang & Agrawala, 2025).
Flow-Based Experts.
We evaluate our method and several monolithic models, including image-to-video (Image2V), depth-to-video (Depth2V), and trajectory-to-video (Traj2V), all using the same Wan2.1 (14B) model as backbone, on a dataset consisting of 3 simulation inputs paired with 4 input images each, giving a total of 12 scenes with examples in Fig. 3. To obtain the initial frame for all methods, we edit the input image with our method described in Section 4.1 to insert the simulated object. Depth maps required as method inputs are rendered from the physical simulator, and input trajectory maps are computed using object centroids. Metrics include the ones from Section 4.1 that are relevant to this task, with additional ones adopted from a video benchmark VBench (Huang et al., 2024): motion smoothness scores introduced in VBench, MUSIQ (Ke et al., 2021) for imaging quality, and ViCLIP (Wang et al., 2023) for text prompts alignment. We further compute foreground motion trajectory accuracy with IoU on SAM2 (Ravi et al., 2024) detected from output videos.
Results are in Table 2 (top half). Depth-to-video and trajectory-to-video alone fail to capture rich non-foreground-object motions, such as cars moving and birds running in Fig. 4. In particular, they tend to compensate for the object’s downward motion with upward camera motions, as visualized by the tracking trajectories in Fig. 4. Image-to-video model does not follow the input motion.
Autoregressive Experts.
We implement the autoregressive construction for Eq. 3 using two generative experts: a next-frame-section video prediction model, FramePack (Zhang & Agrawala, 2025), and a Gaussian distribution $p^{\text{sim}}(x)$ centered on physics simulator renderings. Denote a video sequence of the physics simulator’s RGB rendering overlayed with the input initial frame as $c^{\text{sim}}$, then $p^{\text{sim}}(x):\propto\exp{(-w\|x-c^{\text{sim}}\|_{2}^{2})}$ with constant $w\in\mathbb{R}$. In this case, $x_{1:t}$ represents the first T+1tT+1-t frame sections of a video $x$. We approximate sampling from this distribution with its MAP solution, which amounts to gradient updates w.r.t. loss $\|x-c^{\text{sim}}\|_{2}^{2}$. Qualitative results are in Fig. 5, suggesting that composing per-expert constraints enforces desired object motion on top of output photorealism.
Full-Scene Simulation.
In the above setting, simulation inputs only describe foreground objects. We further evaluate cases where full-scene simulation is available. We use the simulation results from 9 released scenes from PhysGen3D (Chen et al., 2025), a method that reconstructs and animates images in physical simulators, as evaluation inputs. To better preserve scene content, we use the flow-inverted noise vector for particle initialization as opposed to random noise, and compare with this inversion baseline. Results are included in Table 2 with qualitative samples in Fig. 7.
Figure 5: Physical-Simulation-Instructed Video Generation compared with the baseline with the same backbone. Our method better adheres to the input object motion trajectory.
Method mIoU ($\uparrow$) VQAScore ($\uparrow$) VQAScore-R ($\uparrow$)
[…]
$L=32$ $\mathbf{0.728}$$\pm 0.170$ $\mathbf{0.904}$$\pm 0.092$ $\mathbf{0.881}$$\pm 0.115$
Table 3: Text-to-Image Generation with Layout Controls on MIG-Bench (Zhou et al., 2024) dataset, comparing our method with varying compute budgets and an application-specific baseline.
4.3 Text-to-Image Generation with Layout Control
Task and Evaluation.
This task aims to generate an image given an input global text prompt and a set of object bounding boxes with paired object text descriptions. We evaluate on 50 scenes randomly sampled from the full MIG-Bench (Zhou et al., 2024) dataset of 800 scenes. Metrics include bounding box mIoU, detected using GroundingDINO (Liu et al., 2024b), and VQAScore (Lin et al., 2024) measuring output alignment with global and regional prompts.
@@ Results.
Tables 3 and 8 contains results on ablating the number of particles $L$ and also comparisons with 3DIS-FLUX (Zhou et al., 2025), a state-of-the-art method designed for text-to-image generation with the same backbone model as ours. In the case of $L=1$, the method reduces to Du et al. (2023). The performance of our method improves with higher computation budgets, and alleviates the need for application-specific designs, such as cross-attention layer intervention as used in Zhou et al. (2025). The ablation on SMC resampling is deferred to Section D.4.
@@ 5 Conclusion
We have introduced a probabilistic framework where knowledge from heterogeneous sources is composed in the form of product distributions, and proposed an efficient sampling algorithm interleaving annealed MCMC sampling and SMC resampling to integrate guidance from multiple generative and discriminative models. Our training-free approach can effectively integrate visual generative priors, VLM guidance, and physics-based constraints. Empirical evaluation suggests that this method generates images and videos with improved controllability and fidelity compared to prior works, providing a practical recipe for controllable visual synthesis tasks.
@@ Appendix A Overview
The appendix contains discussions (Appendix B), extended descriptions of the framework (Appendix C), and additional experimental results (Appendix D) and details (Appendix E).
@@ Appendix B Discussions
All image experiments are run using 1 NVIDIA H200 GPUs and all video experiments with 2 NVIDIA H200 GPUs. Generating one sample in Section 4.1 takes approximately 4min. It takes 5min for flow-based and 30min for autoregressive-based models in Section 4.2. Wall-clock time for the task in Section 4.3 with different numbers of particles is included in Table 4.
@@
The proposed method requires extra computation due to intermediate MCMC steps and parallel sampling compared to vanilla feedforward inference of the image or video generative model backbones. Furthermore, the sampling efficacy relies on the assumption that expert predictions are reasonably compatible, and strong incompatibilities might require other sampling techniques or training. This work proposes a framework and provides empirical validation with small number of experts and we leave large-scale sampling as future work.
@@
Appendix C Framework Details
C.1 Flow Models
We use bar notations for pre-trained models in the following discussions.
Consider the $i$-th generative expert, which is a flow model that defines a time-dependent continuous probability path $\bar{p}^{(i)}(x;\bar{t}),x\in\mathbb{R}^{d},\bar{t}\in\left[0,1\right]$, with boundary conditions $\bar{p}^{(i)}(x;0)=\mathcal{N}(0,I)$ and $\bar{p}^{(i)}(x;1)\approx p_{\text{data}}^{(i)}(x)$. Each flow model is associated with a scheduler $\bar{\alpha},\bar{\sigma}:\left[0,1\right]\rightarrow\mathbb{R}$ that defines a conditional probability path $\bar{p}^{(i)}(x\mid y;\bar{t})=\mathcal{N}(x\mid\bar{\alpha}(\bar{t})y,\bar{\sigma}(\bar{t})I),y\sim p_{\text{data}}^{(i)}(y)$, and is trained to predict a velocity field $\bar{v}^{(i)}(x;\bar{t},\theta_{i})$ that generates the marginalized path $\bar{p}^{(i)}(x;\bar{t})$, where $\theta_{i}$ denotes model parameters. While the training-time scheduler $(\bar{\alpha},\bar{\sigma})$ can differ across experts, they can be aligned during inference time via a scale-time transformation (Lipman et al., 2024).